1a) Using the double declining depreciation method for an asset with a useful life of 8 years. What is the depreciation expense for year 2 on an asset costing $100,000?

2a) You purchase a home and secure a 30 year equal payment loan for $200,000 at a interest rate of 5.25% APR compounded monthly. After 5 years the interest rate drops to 4.75% APR compounded monthly. The bank is charging 2 points to originate the new loan.

How many months do you need to stay in the house after the refinance to make the refinance a benefit (Round to next month)?
Answer #1

1.

Calculation of depreciation expense for the year 2:

Double-declining depreciation rate = 2 * (1/8) = 0.25 or 25%
Depreciation for the year 1 = $100,000 × 25% = $25,000
Book value at the end of the first year = $100,000 - $25,000 = $75,000

Depreciation for year 2 = $75,000 × 25% = $18,750.

Depreciation expense for the year 2 is $18,750.

Note: As per HOMEWORKLIB RULES, only first question can be answered. post the rest of the questions separately.